OpenStudy (anonymous):

Kelly went to a hockey game with her family. They bought 4 bags of cotton candy and 2 souvenir cups and spent $36. Marissa’s family went to the same game, bought 7 bags of cotton candy and 3 souvenir cups, and spent $59. How much did one souvenir cup cost at the game?

OpenStudy (pfenn1):

Let x = cost of cotton candy and y = cost of souvenir cups The total spent equals x times # of cotton candy bought and y times # cups bought. Can you write that equation for Kelly's family?

OpenStudy (anonymous):

$7 $8 $6 $5

OpenStudy (anonymous):

those are the answer choices

OpenStudy (anonymous):

it is 8

OpenStudy (pfenn1):

@xXDWEAHXx87, we are supposed to help them figure their problems out, not just give them the answer. The idea is to help them learn, not just finish their homework.

OpenStudy (anonymous):

o sorry ill explain

OpenStudy (anonymous):

set it as a system of equations: 4x+2y=36 7x+3y=59 divide the whole top equation by two to simplify it: 2x+y=18, now multiply that whole equation by negative 3 so when you add the two equations, the y's cancel out: -6x-3y=-54 now line up the two equations and add: -6x-3y=-54 7x+3y=59 x+0y=5 x=5 now plug that in into either equation: 4(5)+2y=36 20+2y=36 2y=16 y=8 Hope that helped!